if ABCD is a square of area of 4 sq cm, find the radius of the circle that passes through all the vertices of the square.

Answer by mananth(16949)

square
Area = 4cm^2
side = 2
Use Pythagoras theorem
diagonal = hypotenuse
diagonal ^2= side1^2+side2^2
diagonal^2=2^2+2^2
diagonal^2=8
diagonal = sqrt(8)
the radius of the sircle will be 1/2 the diagonal
sqrt(8)/2
=2sqrt(2)/2
=sqrt(2)
m.ananth@hotmail.ca